OpenStudy (anonymous):

Simplify using order of operations: 3[(3) third power] - (8 x 3). A. -15 B. -3 C. -2 D. 3

OpenStudy (anonymous):

@ClawTalon !!! come to mii cutiee:)

OpenStudy (anonymous):

3^3-(8-3)

OpenStudy (anonymous):

@Itrytohelpu

OpenStudy (anonymous):

3^3= 27 8*3= 24 so.. 27-24=3 ... D is your answer

OpenStudy (anonymous):

3 to the third power is 27, and 8 x 3 = 24. 27 -24 = 3

OpenStudy (anonymous):

thanks